OpenStudy (anonymous):

If you have a rectangular space with dimensions of 200 ft. x 85 ft., and 32,500 square feet is required by 65 cars, do you have sufficient space for 86 cars? ____ yes ___no

OpenStudy (anonymous):

@radar can you help me with this please

OpenStudy (anonymous):

200x85=17,000 17,000/32,500x65=35

OpenStudy (radar):

Step 1. Calculate the area you have. 200 X 85 =17000 sq ft. Step 2. Calculate the area required to park one car 32500/65 = 500 sq ft Step 3. find amount required to park 86 cars, 86 X 500 =43000

OpenStudy (radar):

Nope you can't park 86 cars, and I believe your method gave the number of cars you could park.

OpenStudy (radar):

Out of curosity, you could park 17000/500 cars = 34 cars.

OpenStudy (anonymous):

I got this one thank you

OpenStudy (anonymous):

Can you help me with another problem if i send it to you

OpenStudy (anonymous):

I have to dowload

OpenStudy (radar):

Yes, or you could post it and others may show alternate methods.

OpenStudy (anonymous):

I dont know how to pdkf

OpenStudy (anonymous):

sorry computer asc acting weird

OpenStudy (radar):

neither do I!

OpenStudy (anonymous):

ill send it over here

OpenStudy (radar):

OK

OpenStudy (radar):

Got it, I will be back, got to let the dog out.

OpenStudy (radar):

The dimensions for the living room is given in two units (ft and in) you will have to convert to the same unit. I suggest ft. 12' 4" converts to 12.33 ft. 19' 4" converts to 19.33 ft. Area is 12.33 X 19.33 = 238.4 sq ft.

OpenStudy (radar):

Rounded to nearest sq ft would be 238 sq ft.

OpenStudy (anonymous):

how did you know its 12.33

OpenStudy (radar):

4 in is 1/3 foot or .333333333333

OpenStudy (radar):

Gotta go, Vegas is on TV. Good luck.

OpenStudy (anonymous):

ok